A university is interested in promoting graduates of its honors program by establishing that the mean GPA of these graduates exceeds 3.50. A sample of 36 honors students is taken and is found to have a mean GPA equal to 3.4. The population standard deviation is assumed to equal 0.40.

a. Construct the hypotheses H0 and HA

b. What is the p-value if a =0.05?

c. At the 5% significance level, does the university reject the null hypothesis?
Answer #1

Given : µ = 3.50 , σ = 0.40, n=36 , 1588094221335_blob.png = 3.40 , α = 0.05

#a)

Step 1) Establish the hypothesis.

A university is interested in promoting graduates of its honors program by establishing that the mean GPA of these graduates exceeds 3.50.

H0 : µ = 3.50               Vs   H1 : µ > 3.50

Step 2) calculating test statistics.

As sample size is greater than 30, We use z statistic

Z = ​​​​ 1588094221233_blob.png     = 3.40-3.50 0.40/36      = -1.5

#b)

Step 3) P value

As α = 0.05 and right tail test.

From standard normal table P(Z > -1.5) = 1- P(Z < -1.5)   = 1- 0.0668 = 0.9332

For z = -1.5   , P value = 0.9332

(a) Table of Areas to the Left of z Z .00 .01 .02 .03 -3.4 .0003 .0003 .0003 .0003 -3.3.0005 00050005.0004 -3.2 .0007.0007.00

#c)

Step 4) Decision rule:

Since ,P value is 0.9332 > 0.05 ( Level of significance ), we do not reject H0

Step 5) Conclusion: There is not sufficient evidence to reject   H0 . At the 5% significance level, the university do not reject the null hypothesis.